The graph of y = 4sin (3x + 2) + 1 has amplitude of:
Answer by ikleyn(52835)

.

4.      The answer is  4 (four).



When you see a function of the form  


    f(x) = A*sin(bx) + c    or   g(x) = A*cos(bx) + c 


where A, b and c are constants, its amplitude is equal to A, the coefficient at sine or cosine.